# Discrete Audio Tokens: More Than a Survey!

### v2

@@ 3.3 Acoustic Language Models Evaluation
Following the rise of LLMs, researchers have extended the generative auto-regressive framework beyond text to discrete representations of speech (Lakhotia et al., 2021), audio (Borsos et al., 2023b), and music (Copet et al., 2023). This modeling approach has proven highly effective across domains, enabling powerful generation capabilities (Défossez et al., 2024). In this section, we begin by examining unconditional speech generation (Speech Language Models) and text-conditioned generation (text-to-speech (TTS)). We then explore audio generation and finally turn to the music modality.
@@ 3.3.1 Speech Language Modeling
Motivated by Maimon et al. (2025a), each SLM is built upon the Qwen-2.5 architecture (Yang et al., 2024a) ($357$M parameters in total, after removing the text embedding tables) and initialized using TWIST (Hassid et al., 2023). The textual embedding tables are replaced with new audio embedding tables corresponding to the audio codebooks. The models are trained using the standard cross-entropy loss reported in Eq. 13. To ensure all SLMs processed a comparable amount of audio during training, we dynamically adjusted the tokens per batch according to their tokenizer’s frame rate. To accommodate multiple codebooks, the delay pattern from MusicGen (Copet et al., 2023) is applied across all SLM models. The models are trained for a total of $50,000$ optimizer steps, with a context length set to $1024$. The audio target batch size is set to include about $2.9$ hours of speech per backpropagation step. We used the Adam optimizer coupled with a linear learning rate scheduler, applying a $1\%$ warmup ratio (corresponding to $500$ steps). All input samples are fed to the SLMs using a packing strategy by concatenating all samples together until having a sequence of the target length. To ensure fairer evaluations across different tokenizers, the number of codebooks is restricted to a maximum of $8$, promoting better alignment between them. All code was developed using the SpeechBrain toolkit (Ravanelli et al., 2024), and Hugging Face Transformers (Wolf et al., 2019).
@@
We use the publicly available dataset LibriHeavy (Kang et al., 2024) containing $56$k hours of transcribed speech, and the official validation and test sets of LibriSpeech (Panayotov et al., 2015).
@@ 3.3.2 Text-to-Speech
We evaluate all models on all samples in the test split that fall within the length limit. To evaluate the vocal quality of synthesized speech, we use a pretrained UTMOS model (Takaaki et al., 2022). For assessing pronunciation accuracy, we transcribe both the ground-truth and generated utterances using the Whisper Large model (Radford et al., 2023) with greedy decoding. We compute the degraded Word-Error-Rate (dWER) by treating the ASR prediction of the ground-truth audio as the reference instead of the original transcription. We report mean UTMOS scores and micro dWER values; that is, the Levenshtein edit distance computed over the entire dataset. To measure speaker fidelity, we compute the cosine similarity (SpkSim) between X-vectors extracted from the generated and reference audio using the base variant of WavLM (Chen et al., 2022), fine-tuned for speaker verification.
@@ 3.3.3 Audio Generation
Table 12: Comparing performance of audio LMs over different tokenizers. We report FAD, KL divergence, and CLAP score on the AudioCaps test set. We also provide metrics for audio reconstruction. We note that ground truth audio in AudioCaps gets CLAP$=0.311$, proving a topline. For more information about the tokenizers see Section 2.
@@ 3.3.4 Music Generation
[…]

### v3

@@ 3.3 Acoustic Language Models Evaluation
Following the rise of LLMs, researchers have extended the generative auto-regressive framework beyond text to discrete representations of speech (Lakhotia et al., 2021), audio (Borsos et al., 2023b), and music (Copet et al., 2023). Acoustic language models refer to models that learn to represent or generate audio signals (including speech, general audio, and music) using this generative autoregressive framework. This modeling approach has proven highly effective across domains, enabling powerful generation capabilities (Défossez et al., 2024). In this section, we begin by examining unconditional speech generation (Speech Language Models) and text-conditioned generation (text-to-speech (TTS)). We then explore audio generation and finally turn to the music modality.
@@ 3.3.1 Speech Language Modeling
Motivated by Maimon et al. (2025a), each SLM is built upon the Qwen-2.5 architecture (Yang et al., 2024a) (357357M parameters in total, after removing the text embedding tables) and initialized using TWIST (Hassid et al., 2023). The textual embedding tables are replaced with new audio embedding tables corresponding to the audio codebooks. The models are trained using the standard cross-entropy loss reported in Eq. 13. To ensure all SLMs processed a comparable amount of audio during training, we dynamically adjusted the tokens per batch according to their tokenizer’s frame rate. To accommodate multiple codebooks, the delay pattern from MusicGen (Copet et al., 2023) is applied across all SLM models. The models are trained for a total of $50,000$ optimizer steps, with a context length set to $1024$. The audio target batch size is set to include about $2.9$ hours of speech per backpropagation step. We used the Adam optimizer coupled with a linear learning rate scheduler, applying a $1\%$ warmup ratio (corresponding to $500$ steps). All input samples are fed to the SLMs using a packing strategy by concatenating all samples together until having a sequence of the target length. To ensure fairer evaluations across different tokenizers, the number of codebooks is restricted to a maximum of $8$, promoting better alignment between them. All code was developed using the SpeechBrain toolkit (Ravanelli et al., 2024), and Hugging Face Transformers (Wolf et al., 2019).
@@
We use the publicly available dataset LibriHeavy (Kang et al., 2024) containing 5656k hours of transcribed speech, and the official validation and test sets of LibriSpeech (Panayotov et al., 2015).
@@ 3.3.2 Text-to-Speech
We evaluate all models on all samples in the test split that fall within the length limit. To evaluate the speech naturalness of synthesized speech, we use a pretrained UTMOS model (Saeki et al., 2022). For assessing pronunciation accuracy, we transcribe both the ground-truth and generated utterances using the Whisper Large model (Radford et al., 2023) with greedy decoding. We compute the degraded Word-Error-Rate (dWER) by treating the ASR prediction of the ground-truth audio as the reference instead of the original transcription. We report mean UTMOS scores and micro dWER values; that is, the Levenshtein edit distance computed over the entire dataset. To measure speaker fidelity, we compute the cosine similarity (SpkSim) between X-vectors extracted from the generated and reference audio using the base variant of WavLM (Chen et al., 2022), fine-tuned for speaker verification.
@@ 3.3.3 Audio Generation
Table 12: Comparing performance of audio LMs over different tokenizers. We report FAD, KL divergence, and CLAP score on the AudioCaps test set. We also provide metrics for audio reconstruction. We note that ground truth audio in AudioCaps gets CLAP=0.311=0.311, proving a topline. For more information about the tokenizers see Section 2.
@@ 3.3.4 Music Generation
For training, we use the genre-balanced Free Music Archive (FMA) dataset (Defferrard et al., 2017), following the setup of stable-audio-open (Evans et al., 2025). All samples are 30 seconds long, and we follow the official split provided in the dataset repository [^18] https://github.com/mdeff/fma. The training set consists of 84,213 samples, totaling $\sim$702 hours. We combine artist, album, keywords, genres, and titles from the metadata to build the text prompt of the model.
@@ Acknowledgments
Appendix A Additional Evaluation for Reconstructed Audio Quality
Table 17 shows the UTMOS V2 scores for different audio tokenizers.
Table 17: UTMOS V2 scores of audio tokenizers across different bitrates.
Model Bitrate (kbps) UTMOS V2
Enc-SMA-24 1.5 1.40
6.0 2.13
24.0 2.71
DAC-SMA-24 1.5 1.40
6.0 2.46
24.0 3.06
ST-S-16 1.0 1.52
4.0 2.92
Mimi-S-24 1.1 2.12
4.4 2.82
DWavL-S-16 1.0 2.16
3.0 2.33
SQ-SMA-16 3.0 3.19
WT-SMA-24 0.98 2.39
WT-S-24 0.52 3.12
Appendix B Computational Setup
Table 18 shows computaional setting for our experiments. Approximate runtimes are provided per run and may vary slightly between tokenizers depending on token rate, number of token streams, and other factors. For downstream tasks time is reported per run, runtimes range from 2 hours (e.g., Keyword Spotting) to 48 hours (e.g., ASR), depending on the task complexity.
Table 18: Computational settings for our experiments.
Experimental Name Computational Resource Approx. Time
Downstream Evaluation 1×A100 (80GB) 2–48 hrs
Reconstructed Audio Quality 1×A6000 (48G) 24 hrs
Speech Language Modeling 2×A100 (40GB) 48 hrs
Text-to-Speech 1×A100 (80GB) 96 hrs
Audio Generation 2×A100 (80GB) 48 hrs
Music Generation 4×A100 (40GB) 48 hrs
Ablation Studies 2×GH200 (100GB) 48 hrs
Appendix C Dataset
For each experiment, we report the dataset, source, and approximate duration of the corresponding dataset. A detailed summary is provided in Table 19. For more information, please refer to the "Dataset" subsection in each benchmark section.
Table 19: Summary of datasets, their specifications, and approximate number of hours used for all experiments in this survey.
Task Dataset ~Hours Data Link
Reconstruction
Speech LibriSpeech (Korvas et al., 2014) 6 Link
Music MUSDB (Rafii et al., 2017) 10 Link
Audio Audioset (Gemmeke et al., 2017) 56 Link
Downstream
ASR (En) LibriSpeech (Korvas et al., 2014) 1,000 Link
ASR (Welsh) CommonVoice 17.0 (Ardila et al., 2020) 8 Link
ASR (Basque) CommonVoice 17.0 (Ardila et al., 2020) 116 Link
Speaker ID / Verification VoxCeleb1 (Nagrani et al., 2017) 350 Link
Emotion Recognition IEMOCAP (Busso et al., 2008) 7 Link
Keyword Spotting Speech Commands (Warden, 2018) 18 Link
Intent Classification SLURP (Bastianelli et al., 2020) 10 Link
Speech Enhancement VoiceBank (Valentini-Botinhao et al., 2016) 10 Link
Speech Separation Libri2Mix (Cosentino et al., 2020) 400 Link
Music Genre Classification GTZAN (Tzanetakis & Cook, 2002) 8 Link
Music Source Separation MUSDB (Rafii et al., 2017) 10 Link
Sound Event Classification ESC-50 (Piczak, 2015) 2 Link
Audio Separation FUSS (Wisdom et al., 2021) 23 Link
Acoustic LM
Speech Language Modeling LibriHeavy (Kang et al., 2024) 56,000 Link
Text-to-Speech LibriTTS (Zen et al. (2019)) 960 Link
Audio Generation Data Mix (see Sec. 3.3.3 for details) 4050 -
Music Generation FMA (Defferrard et al., 2017) 702 Link
Ablation
Speech Data Mix (see Sec. 4 for details) 1,000 -
Music Data Mix (see Sec. 4 for details) 1,000 -
Audio Data Mix (see Sec. 4 for details) 1,000 -